OpenStudy (anonymous):

The length of a rectangle is 8 in. more than its width. The perimeter is 52 in. What is the length of the rectangle? @pooja195

pooja195 (pooja195):

\[\huge~\rm~L=w+8\] The formula for perimeter is \[\huge~\rm~P=2L+2w\] We know that the perimeter is 52in

OpenStudy (anonymous):

i dont understand

OpenStudy (anonymous):

oh ok

pooja195 (pooja195):

Let's try an easier approach .-. 2L+2W=52 <---equation for perimeter W=L-8 <---re-arranged equation for length

pooja195 (pooja195):

Basically take equation 2 and plug it in place of w in equation 1 \[\huge~\rm~2L+2(L-8)=52 \] From here solve for L

OpenStudy (anonymous):

4L(L-8)=52?

pooja195 (pooja195):

Can you show me the steps you did to get that so i can find your mistake?

OpenStudy (anonymous):

i added 2L+2

OpenStudy (anonymous):

Solve the following for L, the length : 2( (L-8) + L ) = 52

pooja195 (pooja195):

Ok Start by distributing 2(L−8) 2 x L =2L 2 x -8=?

OpenStudy (anonymous):

2x-8= -16?

pooja195 (pooja195):

Good Good so now we have \[\huge~\rm~2L+2L-16=52 \] combine like terms

OpenStudy (anonymous):

this is the part that always gets me

pooja195 (pooja195):

Combine the l's together 2+2=?

OpenStudy (anonymous):

4L

pooja195 (pooja195):

Good \[\huge~\rm~4L-16=52\] now add 16 to both sides 52+16=?

OpenStudy (anonymous):

4L=36

pooja195 (pooja195):

56+12=36?

OpenStudy (anonymous):

?

pooja195 (pooja195):

52+16=?

OpenStudy (anonymous):

68

pooja195 (pooja195):

4L=68 divide both sides by 4 4/4=1 just L 68/4=?

OpenStudy (anonymous):

68/4=17

pooja195 (pooja195):

Good we we have L=17 :)

OpenStudy (anonymous):

WOW Thank you

pooja195 (pooja195):

yw ^_^